# Robust Long-Range Perception Against Sensor Misalignment in Autonomous Vehicles

### v1

@@ 1 Introduction
  • We propose learning approach with synthetic data augmentation for the sensor misalignment prediction task where ground-truth data is not readily available.
  • Our model predicts well calibrated uncertainty values for the corresponding misalignment predictions, which can be used further to construct accurate estimate of misalignment over time.
  • We show, that correcting incoming misaligned data with predicted misalignment helps achieve higher detection performance at longer ranges.
  • We propose a powerful sensor augmentation method that improves representation learning in network and makes it less susceptible to noise and misalignment.
@@ 2.1 Calibration
Traditionally, LiDAR and camera offline calibration can be categorized into target-based and target-less methods. Target-based calibration, relies on reference objects such as checkerboards [36, 11], ArUco [13], and circular grid [6] to forge a correspondence between LiDAR data and camera images. These methods typically leverage distinctive features like planes and corners to optimize the calibration process, providing high precision within controlled environments. However, they are often constrained by the necessity of such constrained settings. In contrast, target-less methods emphasize the use of natural environmental features for calibration. This trend moves away from the need for controlled environments by exploiting scene-based features (e.g., lines [28, 1] and edges [35, 21]). Sensor motion [14] and gradient [34] based methods are also introduced to further reduce the dependence on scene-based features. These methods offer a more flexible approach and can be applied in situations without the need for elaborate setups. However, maintaining desired level of calibration precision is challenging due to varying environmental and operational factors such as temperature changes, road vibrations, and equipment wear and tear, all of which can introduce calibration errors on an initially well-calibrated system.
@@
Therefore, online calibration is needed to better handle these conditions on the road. In [22], authors monitored the calibration by computing alignment between image edges and projected LiDAR points. An objective function is proposed to model the information of discontinuity from the image and the depth from LiDAR points. Additionally, inspired by the recent success of deep neural networks, online learning-based approaches are also introduced to calibration. RegNet [32], the first deep convolutional neural network for LiDAR-camera calibration marks a significant breakthrough in this field. It is trained on a large amount of miscalibration data to accurately estimate extrinsic parameters. Later end-to-end networks like CalibNet [15] introduce a geometric supervision method to increase their precision, and LCCNet [26] exploits a cost volume layer for feature matching, which allows it to express the correlation between different sensors.
@@
Although online calibration methods have been introduced, it’s important to acknowledge that perfect calibration might not be achievable and the perception system must also be robust enough to handle unexpected discrepancies in sensor data. For instance, [15] achieves impressive mean error in yaw orientation estimate. However, it is still high enough that LiDAR points corresponding to a pedestrian at a 400-meter range would be projected laterally onto a different lane. Moreover, these methods are not multi-task methods and require a standalone deep network alongside a detection network increasing the cost for resource-constrained platforms.
@@ 3.2.3 Loss Functions
Below, we iterate over all the loss function used to train the multi-task network mentioned in the previous section.
@@
$\mathcal{L}_{\text{class}}=-\frac{1}{N}\sum_{i}^{N}\alpha_{i}(1-p_{i})^{\gamma}\log(p_{i})$ (4)
@@
$\mathcal{L}_{\text{2d}}$ $=\frac{1}{N}\sum_{i=1}^{N}\Bigg{(}\frac{\|o^{x}_{i}-o^{x*}_{i}\|_{1}}{b_{o^{x}_{i}}}+\frac{\|o^{y}_{i}-o^{y*}_{i}\|_{1}}{b_{o^{y}_{i}}}$ (5)
$\quad+\frac{\|w^{\text{2d}}_{i}-w_{i}^{\text{2d}*}\|_{1}}{b_{w_{i}^{\text{2d}}}}+\frac{\|h^{\text{2d}}_{i}-h_{i}^{\text{2d}*}\|_{1}}{b_{h_{i}^{\text{2d}}}}$
$\quad+\log(b_{o^{x}_{i}}b_{o^{y}_{i}}b_{w_{i}^{\text{2d}}}b_{h_{i}^{\text{2d}}})\Bigg{)}$
@@
$\mathcal{L}_{\text{3d}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\|\hat{o}^{x}_{i}-\hat{o}^{x*}_{i}\|_{1}}{b_{\hat{o}^{x}_{i}}}+\frac{\|\hat{o}^{y}_{i}-\hat{o}^{y*}_{i}\|_{1}}{b_{\hat{o}^{y}_{i}}}\right.$ (6)
$+\left.\frac{\|w^{\text{3d}}_{i}-w_{i}^{\text{3d}*}\|_{1}}{b_{w_{i}^{\text{3d}}}}+\frac{\|l^{\text{3d}}_{i}-l_{i}^{\text{3d}*}\|_{1}}{b_{l_{i}^{\text{3d}}}}+\frac{\|h^{\text{3d}}_{i}-h_{i}^{\text{3d}*}\|_{1}}{b_{h_{i}^{\text{3d}}}}\right.$
$+\left.\log(b_{\hat{o}^{x}_{i}}b_{\hat{o}^{y}_{i}})+\log(b_{w_{i}^{\text{3d}}}b_{h_{i}^{\text{3d}}}b_{l_{i}^{\text{3d}}})\right.$
$+\|d_{i}-d_{i}^{*}\|_{1}+\|\phi_{i}-\phi_{i}^{*}\|_{1}\Bigg{)}$
@@
$\begin{split}\mathcal{L}_{\text{miscal}}&=\frac{\|\theta^{\text{pitch}}-\theta^{\text{pitch*}}\|_{1}}{b_{\theta^{\text{pitch}}}}+\frac{\|\theta^{\text{yaw}}-\theta^{\text{yaw*}}\|_{1}}{b_{\theta^{\text{yaw}}}}\\ &\quad+\frac{\|\theta^{\text{roll}}-\theta^{\text{roll*}}\|_{1}}{b_{\theta^{\text{roll}}}}+\log(b_{\theta^{\text{pitch}}}b_{\theta^{\text{yaw}}}b_{\theta^{\text{roll}}})\end{split}$ (7)
@@ 4 Verification
We employ a fault injection methodology that simulates misalignment conditions by perturbing the extrinsic parameters on a carefully curated test dataset. Specifically, we introduce controlled misalignment by altering the extrinsic parameters of the LiDAR sensor, which defines its spatial orientation and position relative to the vehicle’s coordinate frame. These faults are injected by perturbing the roll, pitch, and yaw angles within a range of -1.0 to 1.0 degrees of rotation. This range is selected to emulate realistic scenarios where minor misalignments can occur due to factors such as vibrations, impacts, or thermal effects.
@@ 4.1.1 Dataset
To validate our proposed approach, we leverage our diverse internal dataset. Our focus in this study is mainly the degradation in long-range detections due to small misalignments. This limits us from using publicly available datasets such as Waymo Dataset which provides 3D labels only until 85 meters range [33] and KITTI [10] with max range of 120 meters. Our internal proprietary dataset provides 3D bounding box labels up to 500m range using Frequency Modulated Continuous Wave (FMCW) LiDAR system. It comprises 43,500 five-second snippets for training and 1,000 for validation, featuring synchronized 8MP camera images with a 30° field of view, LiDAR data from FMCW LiDAR system. More importantly, this dataset enables us to verify our method’s effectiveness on distant objects (upto 500 meter range) under diverse conditions.
From the dataset, we extract 5-second long continuous snippets of sensor data stream. To simulate realistic misalignment scenarios, we employ purpose built fault injector tool to perturb the incoming LiDAR point data within each snippet, introducing roll, pitch and yaw perturbations uniformly sampled in the range of [-1 degree to 1 degree] with an interval of 0.1 degree with total 1000 snippets of data.
@@
Per Frame 0.9813 0.9637 0.0700 0.0313 0.0241
Snippet (w/o uncertainty) 0.9821 0.9644 0.0620 0.0298 0.0184
Snippet (w uncertainty) 0.9861 0.9650 0.0450 0.0290 0.0141
Table 1: Performance metrics per frame, per snippet with fused predictions without considering uncertainty, and per snippet with fused predictions considering uncertainty (filtering out estimates with high predicted uncertainty and performing weighted average over remaining frames).
@@ 4.1.2 Evaluation Metrics
Here we describe different metrics used to demonstrate effectiveness of our approach.
@@
Misalignement Detection Accuracy: The misalignment detection accuracy metric (see MDA columns in Table 1) evaluates our method’s ability to reliably identify sensor alignment degradation while minimizing false alarms. We evaluate result as positive when our method predicts misalignment of $>0.1$ degree and negative when our method predicts misalignment of $<0.1$ degree. True-Positive is considered, when both predicted and injected misalignment are $>0.1$ degree and True-Negative is considered when both predicted and injected misalignment are $<0.1$ degrees. Using this, we report recall and precision across all snippets under varied conditions to comprehensively assess misalignment detection performance.
@@
Misalignment Estimation Error: We evaluate misalignment estimation error at both the per-frame and per-snippet levels. For per-frame evaluation, we directly compare the predicted misalignment values against the injected ground truth misalignment. For per-snippet evaluation, we compute the mean of all frame-level misalignment predictions within a snippet. Additionally, we leverage the predicted uncertainty estimates to fuse the outputs by ignoring predictions with high uncertainty (sigma $>0.3$ degrees) and computing a weighted average over the remaining predictions as described in section 3.3.
@@
Perception Performance Metrics: Additionally, we measure the 3D detection performance of our multi-task model against a baseline and ensure that our module-level mitigation method strictly increases perception performance across all conditions (see Table 2). We employ the max F1 score metric (IoU=0.1) to evaluate the Bird’s Eye View (BEV) object detection performance for actors at different ranges under various fault injection scenarios. As a baseline, we show the result from the model that is not trained with any perturbation. We compare the results of our proposed network architecture with and without applying the correction. This comprehensive evaluation helps us demonstrate the tangible benefits of our mitigation strategy in improving perception robustness and maintaining reliable operation in the presence of sensor misalignment faults.
BEV Max-F1 @ 0.1 IoU for Vehicles $\uparrow$
@@
Table 2: Comparison of 3D vehicle detection performance. The ”Proposed (No Correction Applied)” column represents the performance of our multi-task model when evaluated on miscalibrated (miscal) and calibrated (calib) data without applying any correction. The ”Proposed (Predicted Correction Applied)” column shows the improved performance achieved by applying the extrinsic parameter correction using the misalignment values predicted by our model. For readers, we highlight the highest score in each row with bold characters for calibrated and miscalibrated test data category individually.
@@ 4.2 Results
This section evaluates the performance of our proposed approach and its impact on enhancing safety across various driving scenarios. For system-level failure mitigation, we focus on the Misalignment Detection Accuracy (MDA) metrics, specifically precision and recall. As shown in Table 1, the results demonstrate the high precision and recall achieved by our method in detecting sensor misalignment across different conditions. Furthermore, incorporating uncertainty estimates in the Per Snippet evaluation leads to improved misalignment estimation accuracy, as evident from the lower Mean Estimated Error values.
The table also presents the mean L1 error for the three misalignment dimensions (pitch, roll, and yaw) considered in this work. By combining the table results with the misalignment estimates in the plot (Figure 4) of the injection LiDAR misalignment, we see that the model demonstrates promising performance in estimating yaw misalignment but struggles with roll misalignment. This behavior could be due to the differences in vertical and horizontal resolution for the LiDAR. LiDAR has a slightly higher horizontal resolution, which makes estimating yaw miscalibration easier compared to Pitch. The roll noise is only visible at the very edge of the image, which typically consists of feature-less areas such as the sky or the ground surface closer to the camera (see Figure 1), making it more difficult for the model to observe feature displacement due to misalignment in roll.
Figure 5: Comparison of Baseline and Proposed Method for 3D Detection: The white boxes are ground truth labels, and the red boxes are model predictions. (a) Range view illustrating the presence of LiDAR miscalibration error in yaw. (b) BEV of the baseline 3D detection performance. The white dots represent LiDAR points. (c) BEV of the proposed method. It shows a clear improvement for objects located beyond 250 meters, highlighting its effectiveness in addressing LiDAR miscalibration errors at long range.
Table 2 compares the performance of our multi-task model and a baseline model for vehicle detection on calibrated and miscalibrated data. We observe that the multi-task model significantly improves detection performance compared to the baseline when evaluated on the same miscalibrated data. Furthermore, when applying correction (see qualitative example in Figure 1) using the predicted misalignment values, the performance improves for ranges beyond 200 meters. Notably, the model already performs adequately without the calibration correction, as it is trained on miscalibrated data, making it inherently robust against misalignment errors and thus reducing the impact of the correction mechanism.
@@
Overall, these results demonstrate the effectiveness of our approach in detecting and mitigating sensor misalignment faults, ultimately improving the safety and reliability of autonomous vehicle’s long-range perception systems in various driving conditions.

### v3

@@ 1 Introduction
  • We propose a learning approach with synthetic data augmentation for the sensor misalignment prediction task where ground-truth data is not readily available.
  • We show that learning misalignment as an auxiliary task serves as a valuable enhancement to object detection models.
  • Our model predicts well-calibrated uncertainty values for the corresponding misalignment predictions, which can be used further to construct accurate estimate of misalignment over time.
  • We show that correcting incoming misaligned data with predicted misalignment helps achieve higher detection performance at longer ranges.
@@ 2.1 Calibration
Traditionally, LiDAR and camera offline calibration can be categorized into target-based and target-less methods. Target-based calibration, relies on reference objects such as checkerboards [37, 12], ArUco [14], and circular grid [6] to forge a correspondence between LiDAR data and camera images. These methods typically leverage distinctive features like planes and corners to optimize the calibration process, providing high precision within controlled environments. While these methods provide high precision in controlled environments, they are limited by the need for such specific settings. In contrast, target-less methods leverage natural environmental features, such as lines (e.g., lines [28, 1] and edges [35, 21]). Yuan et al.[36] extracted more accurate and reliable LiDAR edges by using voxel cutting and plane fitting techniques. However, maintaining a desired level of calibration precision is challenging due to varying environmental and operational factors such as temperature changes, road vibrations, and equipment wear and tear, all of which can introduce calibration errors on an initially well-calibrated system.
@@
Therefore, online calibration is needed to better handle these conditions on the road. In [22], authors monitored the calibration by computing alignment between image edges and projected LiDAR points. An objective function is proposed to model the information of discontinuity from the image and the depth from LiDAR points. Additionally, inspired by the recent success of deep neural networks, online learning-based approaches are also introduced to calibration. RegNet [32], the first deep convolutional neural network for LiDAR-camera calibration marks a significant breakthrough in this field. It is trained on a large amount of miscalibration data to accurately estimate extrinsic parameters. Zhu et al. [38] proposed a semantic segmentation method that first extracts semantic features and combines non-monotonic line search and subgradient ascent to estimate optimal calibration parameters. Sun et al. [33] introduced an instance segmentation approach with a cost function that accounts for the matching degree based on the appearance and centroids of key targets in point cloud and image pairs. End-to-end networks, such as CalibNet [15] incorporate geometric supervision to enhance precision, while LCCNet [26] exploits a cost volume layer for feature matching, enabling it to capture correlations between different sensors.
@@
Although online calibration methods have been introduced, it’s important to acknowledge that perfect calibration might not be achievable and the perception system must also be robust enough to handle unexpected discrepancies in sensor data. For instance, [15] achieves impressive mean error in yaw orientation estimate. However, it is still high enough that LiDAR points corresponding to a pedestrian at a 400-meter range would be projected laterally onto a different lane. Moreover, these methods are not multi-task methods and require a standalone network alongside a detection network increasing the cost for resource-constrained platforms.
@@ 3.2.3 Loss Functions
Below, we iterate over all the loss functions used to train the multi-task network mentioned in the previous section. We apply different weights $W_{\{c,o,s,\hat{o},\hat{s},d,\phi,\theta\}}$ on each loss for multi-task learning.
@@
$\mathcal{L}_{\text{class}}=-\frac{W_{c}}{N}\sum_{i}^{N}\alpha_{i}(1-p_{i})^{\gamma}\log(p_{i})$ (4)
@@
$\mathcal{L}_{\text{2D}}$ $=\frac{1}{N}\sum_{i=1}^{N}\Bigg{(}W_{o}(\frac{\|o^{x}_{i}-o^{x*}_{i}\|_{1}}{b_{o^{x}_{i}}}+\frac{\|o^{y}_{i}-o^{y*}_{i}\|_{1}}{b_{o^{y}_{i}}})$ (5)
$\quad+W_{s}(\frac{\|w^{\text{2D}}_{i}-w_{i}^{\text{2D}*}\|_{1}}{b_{w_{i}^{\text{2D}}}}+\frac{\|h^{\text{2D}}_{i}-h_{i}^{\text{2D}*}\|_{1}}{b_{h_{i}^{\text{2D}}}})$
$\quad+W_{o}\log(b_{o^{x}_{i}}b_{o^{y}_{i}})+W_{s}\log(b_{w_{i}^{\text{2D}}}b_{h_{i}^{\text{2D}}})\Bigg{)}$
@@
$\mathcal{L}_{\text{3D}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(W_{\hat{o}}(\frac{\|\hat{o}^{x}_{i}-\hat{o}^{x*}_{i}\|_{1}}{b_{\hat{o}^{x}_{i}}}+\frac{\|\hat{o}^{y}_{i}-\hat{o}^{y*}_{i}\|_{1}}{b_{\hat{o}^{y}_{i}}})\right.$ (6)
$+\left.W_{\hat{s}}(\frac{\|w^{\text{3D}}_{i}-w_{i}^{\text{3D}*}\|_{1}}{b_{w_{i}^{\text{3D}}}}+\frac{\|l^{\text{3D}}_{i}-l_{i}^{\text{3D}*}\|_{1}}{b_{l_{i}^{\text{3D}}}}+\frac{\|h^{\text{3D}}_{i}-h_{i}^{\text{3D}*}\|_{1}}{b_{h_{i}^{\text{3D}}}})\right.$
$+\left.W_{\hat{o}}\log(b_{\hat{o}^{x}_{i}}b_{\hat{o}^{y}_{i}})+W_{\hat{s}}\log(b_{w_{i}^{\text{3D}}}b_{h_{i}^{\text{3D}}}b_{l_{i}^{\text{3D}}})\right.$
$+W_{d}\|d_{i}-d_{i}^{*}\|_{1}+W_{\phi}\|\phi_{i}-\phi_{i}^{*}\|_{1}\Bigg{)}$
@@
$\begin{split}\mathcal{L}_{\text{miscal}}&=W_{\theta}\Bigg{(}\frac{\|\theta^{\text{pitch}}-\theta^{\text{pitch*}}\|_{1}}{b_{\theta^{\text{pitch}}}}+\frac{\|\theta^{\text{yaw}}-\theta^{\text{yaw*}}\|_{1}}{b_{\theta^{\text{yaw}}}}\\ &\quad+\frac{\|\theta^{\text{roll}}-\theta^{\text{roll*}}\|_{1}}{b_{\theta^{\text{roll}}}}+\log(b_{\theta^{\text{pitch}}}b_{\theta^{\text{yaw}}}b_{\theta^{\text{roll}}})\Bigg{)}\end{split}$ (7)
@@
We choose multi-task weights $W_{o}$ = $1.0,W_{c}$ = $2.0,W_{s}$ = $0.1,W_{\hat{o}}$ = $0.25,W_{\hat{s}}$ = $1.0,W_{d}$ = $1.5,W_{\phi}$ = $0.1,$ and $W_{\theta}$ = $0.4$ for the experiments.
@@ 4 Verification
We employ a fault injection methodology that simulates misalignment conditions by perturbing the extrinsic parameters on a carefully curated test dataset. Specifically, we introduce controlled misalignment by altering the extrinsic parameters of the LiDAR sensor, which defines its spatial orientation and position relative to the vehicle’s coordinate frame. These faults are injected by perturbing the roll, pitch, and yaw angles within a range of -1.0 to 1.0 degrees of rotation. This range is selected to emulate realistic scenarios where minor misalignments can occur due to factors such as vibrations, impacts, or thermal effects. In the experiment section, we show such minor misalignment is enough to have a severe impact on long range perception (45% regression in 300-400 meters from Table 3).
@@ 4.1.1 Dataset
To validate our proposed approach, we leverage our diverse internal dataset. Our focus in this study is mainly the degradation in long-range detections due to small misalignments. This limits us from using publicly available datasets such as Waymo Dataset which provides 3D labels only until 85 meters range [34] and KITTI [11] around 70 meters. To provide better reproducibility and to show applicability of this approach for short range perception, we also conduct experiments on Waymo dataset and compare the results against available approaches. The limitations will be further discussed in the following sections.
Our internal proprietary dataset provides 3D bounding box labels up to 500m range using Frequency Modulated Continuous Wave (FMCW) LiDAR system. It comprises 43,500 five-second snippets for training and 1,000 for validation, featuring synchronized 8MP camera images with a 30° field of view, LiDAR data from FMCW LiDAR system. More importantly, this dataset enables us to verify our method’s effectiveness on distant objects (upto 500 meter range) under diverse conditions. From the dataset, we extract 5-second long continuous snippets of sensor data stream. To simulate realistic misalignment scenarios, we employ purpose built fault injector tool to perturb the incoming LiDAR point data within each snippet, introducing roll, pitch and yaw perturbations uniformly sampled in the range of [-1 degree to 1 degree] with an interval of 0.1 degree with total 1000 snippets of data.
On Internal Proprietary Long Range Dataset with $\delta\in[-1^{\circ},1^{\circ}]$
@@
Per Frame 0.9813 0.9637 $0$$\pm 0.017$ 0.0313 ±$\pm 0.008$ 0.0241 ±$\pm 0.006$
Snippet (w/o uncertainty) 0.9821 0.9644 0.0620 ±$\pm 0.015$ 0.0298 ±$\pm 0.005$ 0.0184 ±$\pm 0.006$
Snippet (w uncertainty) 0.9861 0.9650 0.0450 ±$\pm 0.008$ 0.0290 ±$\pm 0.001$ 0.0141 ±$\pm 0.0007$
Table 1: Performance metrics with CenterNet baseline [7] per frame, per snippet with fused predictions without considering uncertainty, and per snippet with fused predictions considering uncertainty (filtering out estimates with high predicted uncertainty and performing weighted average over remaining frames).
@@ 4.1.2 Evaluation Metrics
Here we describe different metrics used to demonstrate the effectiveness of our approach. For the internal long-range dataset, we apply a 1.0 degree LiDAR perturbation, as previously described. For the Waymo dataset, a 5.0 degree perturbation is applied to create more pronounced misalignment in short-range scenarios.
Method Pitch° Yaw° Roll°
Yuan [36] 0.411 0.325 0.423
Zhu [38] 0.421 0.301 0.305
Sun [33] 0.202 0.211 0.171
Ours 0.248 0.068 0.309
Table 2: Comparison of different methods for pitch, yaw, and roll error on Waymo dataset [34].
@@
Misalignment Detection Accuracy: The misalignment detection accuracy metric (see MDA columns in Table 1) evaluates our method’s ability to reliably identify sensor alignment degradation while minimizing false alarms. We evaluate results as positive when our method predicts misalignment of $>0.1$ degree and negative when our method predicts misalignment of $<0.1$ degree. True-Positive is considered, when both predicted and injected misalignment are $>0.1$ degree and True-Negative is considered when both predicted and injected misalignment are $<0.1$ degrees. Using this, we report recall and precision across all snippets under varied conditions to comprehensively assess misalignment detection performance.
@@
Misalignment Estimation Error Analysis: We evaluate the misalignment estimation error at both the per-frame and per-snippet levels. For per-frame evaluation, we directly compare the predicted misalignment values against the injected ground truth misalignment. For per-snippet evaluation, we compute the mean of all frame-level misalignment predictions within a snippet. Additionally, we leverage the predicted uncertainty estimates to fuse the outputs by ignoring predictions with high uncertainty (sigma $>0.3$ degrees) and computing a weighted average over the remaining predictions as described in section 3.3.
Misalignment Estimation Error Benchmark: To better evaluate the effectiveness of our approach, we compare it with other target-less learning methods on the Waymo dataset [34]. We leverage the experiment settings from Sun et al. [33], where perturbations are applied differently at each group, and the mean average error is calculated as the final result. However, we compute the average over 30,000 frames, as opposed to the 60 frames in the original setting [33], since we observed result fluctuations across different groups. The comparison is presented in Table 2.
@@
Perception Performance Metrics: In addition to CenterNet, we include SpotNet [9] in this experiment to demonstrate the effectiveness of the proposed method as an auxiliary task. SpotNet is a long range, image-centric, and LiDAR-anchored approach that jointly learns 2D and 3D detection tasks. We measure the 3D detection performance of our multi-task model against a baseline and ensure that our module-level mitigation method strictly increases perception performance across all conditions (see Table 3). We employ the max F1 score metric (IoU=0.1) to evaluate the Bird’s Eye View (BEV) object detection performance for actors at different ranges under various fault injection scenarios. As a baseline, we show the result from the model that is not trained with any perturbation. We compare the results of our proposed network architecture with and without applying the correction. This comprehensive evaluation helps us demonstrate the tangible benefits of our mitigation strategy in improving perception robustness and maintaining reliable operation in the presence of sensor misalignment faults.
Additionally, for the Waymo dataset, we include only the results from SpotNet due to a label association issue. In Waymo, the 2D and 3D labels are not linked, preventing us from directly associating 2D labels with their corresponding 3D labels. As a result, we are unable to train CenterNet on 3D detection, since our method relies on 2D labels to encode target image pixels and then trains the 3D attributes from the associated 3D labels. Therefore, we compare only the 2D detection task on the Waymo dataset (see Table 4).
BEV Max-F1 @ 0.1 IoU for Vehicles with CenterNet [7] $\uparrow$
@@
*BEV Max-F1 @ 0.1 IoU for Vehicles with SpotNet [9] $\uparrow$
calib miscal calib miscal calib miscal
200-300 0.804 0.360 0.794 (-1.2%) 0.789 (+119%) 0.789 (-1.8%) 0.792 (+124.3%)
300-400 0.734 0.250 0.726 (-1.0%) 0.718 (+187%) 0.726 (-1.0%) 0.721 (+188%)
400-500 0.640 0.220 0.620 (-3.12%) 0.601 (+173%) 0.626 (-3.06%) 0.626 (+184%)
Table 3: Comparison of 3D vehicle detection performance. The ”Proposed (No Correction Applied)” column represents the performance of our multi-task model when evaluated on miscalibrated (miscal) and calibrated (calib) data without applying any correction. The ”Proposed (Predicted Correction Applied)” column shows the improved performance achieved by applying the extrinsic parameter correction using the misalignment values predicted by our model. For readers, we highlight the highest score in each row with bold characters for calibrated and miscalibrated test data category individually. *The dataset count slightly differs.
@@ 4.2 Results
This section evaluates the performance of our proposed approach and its impact on enhancing safety across various driving scenarios. For system-level failure mitigation, we focus on the Misalignment Detection Accuracy (MDA) metrics, specifically precision and recall. As shown in Table 1, the results demonstrate the high precision and recall achieved by our method in detecting sensor misalignment across different conditions. Additionally, incorporating uncertainty estimates in the Per Snippet evaluation improves misalignment estimation accuracy, as evidenced by the lower Mean Estimated Error values.
Table 2 presents results that are comparable to the benchmark in pitch and roll, while outperforming in yaw. By combining these table results with the misalignment estimates shown in the plot (Figure 4) of the injection LiDAR misalignment, we see that the model demonstrates promising performance in estimating yaw misalignment but struggles with roll misalignment. This behavior could be due to the differences in vertical and horizontal resolution for the LiDAR. LiDAR has a slightly higher horizontal resolution, which makes estimating yaw miscalibration easier compared to Pitch. The roll noise is only visible at the very edge of the image, which typically consists of feature-less areas such as the sky or the ground surface closer to the camera (see Figure 1), making it more difficult for the model to observe feature displacement.
Table 3 compares the performance of our multi-task model and a baseline model for vehicle detection on calibrated and miscalibrated data. We observe that the multi-task model significantly improves detection performance compared to the baseline when evaluated on the same miscalibrated data. It is important to note the significant performance drop in SpotNet [9] under miscalibrated scenarios. This is because SpotNet is a LiDAR-anchored approach, making it highly sensitive to miscalibration. When there is misalignment between the projected sensor points and the image, the model mismatches the sensor data with the image background features, leading to decreased accuracy. Furthermore, when applying correction (see qualitative example in Figure 1) using the predicted misalignment values, the performance improves for ranges beyond 200 meters. Notably, the model already performs adequately without the calibration correction, as it is trained on miscalibrated data, making it inherently robust against misalignment errors and thus reducing the impact of the correction mechanism. Table 4 also shows that the proposed method improves short range 2D detection on the Waymo dataset
@@
On Short Range Waymo Dataset[34] with SpotNet [9] with $\delta\in[-5^{\circ},5^{\circ}]$
Approach Eval Data Max-F1 for 2D Detection
Baseline Calibrated 0.693
MisCalibrated 0.640
Proposed w/o input correction Calibrated 0.712
MisCalibrated 0.698
Table 4: 2D detection performance metrics shows improvement on short-range perception with our proposed augmentation learning approach.(note: For short-range waymo dataset, we inject $\delta\in[-5,5]$ degree of miscalibration)